OpenStudy (anonymous):

An object was dropped without initial speed from 7 m above thr ground . the speed object when it reached 2 m from the ground was....

OpenStudy (anonymous):

v^2=u^2 + 2aS v= 0+square_root(2*g*5) = 10 if you take g=10 or root(98) if you take it = 9.8